Mr and Mrs Wombat live in a house in which the floor of every room is a square and covered with identical tiles. Mrs Wombat's room contains 101 tiles more than Mr Wombat's room. How many tiles does Mrs Wombat's room contain?
Answer by AnlytcPhil(1810)

Mr and Mrs Wombat live in a house in which the floor of 
every room is a square and covered with identical tiles. 
Mrs Wombat's room contains 101 tiles more than Mr Wombat's 
room. How many tiles does Mrs Wombat's room contain?

Let x = number of tiles in each side of Mr. Wombat's room
Let y = number of tiles in each side of Mrs. Wombat's room

Then there are x² tiles in Mrs. Wombat's room.
And there are y² tiles in Mr. Wombat's room.

>>...Mrs. Wombat's room contains 101 tiles more than Mr. 
Wombat's room...<<

             y² = x² + 101

        y² - x² = 101

 (y - x)(y + x) = 101

The two parentheses on the left must be a factorization
of 101.  Since 101 is a prime number, its only 
factorization is  1×101. Since (y - x) is smaller than
(y + x), we must have

           y - x = 1
and 
           y + x = 101

So we have this system

           y - x = 1
           y + x = 101

Solving that gives x = 51 and y = 50 

So Mrs. Wombat's room has 51 tiles on each side and
Mr. Wombat's room has 50 tiles on each side.

The question is:

>>...How many tiles does Mrs Wombat's room contain?...<<

So her room has 51×51 or 2601 tiles

Checking:
Mrs. Wombat's room has 2601 tiles
Mr. Wombat's room has 50×50 of 2500 tiles
Since 2601 is 101 more than 2500, the solution is correct.

Edwin McCravy
AnlytcPhil@aol.com